OpenStudy (anonymous):

The volume of a square pyramid whose base side is 9 cm and height is 5 cm is _____ cm³.

OpenStudy (anonymous):

Volume = (1/3)*(base)^2*height plug the values into the formula Volume=(1/3)*(9)(9)*5 =(9)(3)*5 =135 cm^2 therefore, Volume=135 cm^2